In 100 seconds a smoke alarm emits 1000000 sound waves. What is the frequency of the sound waves
1
Expert's answer
2015-09-18T09:06:11-0400
The frequency of the sound waves equals the number of vibrations produced for 1 second.
Thus, ν = 1000000/100 s = 10000 s-1 = 10000 Hz = 10 kHz